# Forgetting-MarI: LLM Unlearning via Marginal Information Regularization

### v1

@@ D.2 Supplementary General Model Capacity Test Results
Table 7 illustrates comprehensive evaluation results across multiple benchmark test.
@@
Task Metric GPT-2 Baseline Unlearn Baseline F-MarI
ARC-Easy acc 0.53 ± 0.01 0.46 ± 0.01 0.48 ± 0.01 0.45 ± 0.01
acc_norm 0.47 ± 0.01 0.42 ± 0.01 0.43 ± 0.01 0.43 ± 0.01
ARC-Challenge acc 0.22 ± 0.01 0.23 ± 0.01 0.22 ± 0.01 0.22 ± 0.01
acc_norm 0.25 ± 0.01 0.27 ± 0.01 0.27 ± 0.01 0.26 ± 0.01
PIQA acc 0.70 ± 0.01 0.66 ± 0.01 0.66 ± 0.01 0.66 ± 0.01
acc_norm 0.69 ± 0.01 0.65 ± 0.01 0.65 ± 0.01 0.66 ± 0.01
Hellaswag acc 0.36 ± 0.00 0.36 ± 0.00 0.36 ± 0.00 0.35 ± 0.00
acc_norm 0.45 ± 0.00 0.43 ± 0.00 0.43 ± 0.00 0.42 ± 0.00
MMLU acc 0.23 ±0.00 0.23 ±0.00 0.24 ±0.00 0.23 ±0.00
- humanities acc 0.25 ±0.01 0.24 ±0.01 0.25 ±0.01 0.24 ±0.01
- other acc 0.24 ±0.01 0.24 ±0.01 0.24 ±0.01 0.24 ±0.01
- social sciences acc 0.22 ±0.01 0.22 ±0.01 0.22 ±0.01 0.22 ±0.01
- stem acc 0.22 ±0.01 0.22 ±0.01 0.23 ±0.01 0.22 ±0.01
Task Metric GPT-2 Baseline Unlearn Baseline F-MarI
WikiText bits/byte 0.841 0.925 0.905 0.910
byte-pplx 1.792 1.898 1.873 1.879
word-pplx 22.61 30.80 28.66 29.19
Table 7: Comprehensive evaluation results across multiple benchmarks for the 10/90 split test with GPT2-LG baselines. When the true underlying stopping criteria is unknown, we choose to stop when the detector fails. This corresponds to F-Mari with lambda = 0.95 and epoch = 10.
@@ E.2 Multiple Detection Test Results
Here, we provide more detailed results in addition to Figure 6 in Section $4.$

### v2

@@ D.5 Supplementary General Model Capacity Test Results
Table 5 and 6 summarize comprehensive evaluation results across multiple benchmark tests.
Task Metric GPT2-LG Baseline Unlearn Baseline F-MarI KL-GA GA GD DPO
ARC-Easy acc 0.53 ± 0.01 0.46 ± 0.01 0.48 ± 0.01 0.46 ± 0.01 0.46 ± 0.01 0.24 ± 0.01 0.45 ± 0.01 0.46 ± 0.01
acc_norm 0.47 ± 0.01 0.42 ± 0.01 0.43 ± 0.01 0.43 ± 0.01 0.43 ± 0.01 0.25 ± 0.01 0.43 ± 0.01 0.43 ± 0.01
ARC-Challenge acc 0.22 ± 0.01 0.23 ± 0.01 0.22 ± 0.01 0.23 ± 0.01 0.24 ± 0.01 0.22 ± 0.01 0.23 ± 0.01 0.23 ± 0.01
acc_norm 0.25 ± 0.01 0.27 ± 0.01 0.27 ± 0.01 0.27 ± 0.01 0.28 ± 0.01 0.26 ± 0.01 0.27 ± 0.01 0.27 ± 0.01
PIQA acc 0.70 ± 0.01 0.66 ± 0.01 0.66 ± 0.01 0.66 ± 0.01 0.66 ± 0.01 0.53 ± 0.01 0.65 ± 0.01 0.65 ± 0.01
acc_norm 0.69 ± 0.01 0.65 ± 0.01 0.65 ± 0.01 0.66 ± 0.01 0.66 ± 0.01 0.51 ± 0.01 0.64 ± 0.01 0.65 ± 0.01
Hellaswag acc 0.36 ± 0.00 0.36 ± 0.00 0.36 ± 0.00 0.35 ± 0.00 0.36 ± 0.00 0.25 ± 0.00 0.35 ± 0.00 0.36 ± 0.00
acc_norm 0.45 ± 0.00 0.43 ± 0.00 0.43 ± 0.00 0.42 ± 0.00 0.42 ± 0.00 0.26 ± 0.00 0.42 ± 0.00 0.42 ± 0.00
MMLU acc 0.23 ±0.00 0.23 ±0.00 0.24 ±0.00 0.23 ±0.00 0.23 ±0.00 0.25 ±0.00 0.23 ±0.00 0.23 ±0.00
- humanities acc 0.25 ±0.01 0.24 ±0.01 0.25 ±0.01 0.24 ±0.01 0.24 ±0.01 0.25 ±0.01 0.24 ±0.01 0.25 ±0.01
- other acc 0.24 ±0.01 0.24 ±0.01 0.24 ±0.01 0.24 ±0.01 0.24 ±0.01 0.27 ±0.01 0.25 ±0.01 0.24 ±0.01
- social sciences acc 0.22 ±0.01 0.22 ±0.01 0.22 ±0.01 0.22 ±0.01 0.22 ±0.01 0.23 ±0.01 0.22 ±0.01 0.22 ±0.01
- stem acc 0.22 ±0.01 0.22 ±0.01 0.23 ±0.01 0.22 ±0.01 0.21 ±0.01 0.24 ±0.01 0.22 ±0.01 0.22 ±0.01
Task Metric GPT2-LG Baseline Unlearn Baseline F-MarI KL-GA GA GD DPO
WikiText bits/byte 0.841 0.925 0.905 0.917 0.929 26.895 0.961 0.939
byte-pplx 1.792 1.898 1.873 1.888 1.904 1.248e+08 1.946 1.917
word-pplx 22.612 30.797 28.662 29.886 31.309 1.972e+43 35.214 32.428
Table 5: Comprehensive evaluation results across multiple benchmarks for the GPT2-LG baselines. WikiText is based on perplexity, so lower is better. The higher score is better for all other tests: ARC, PIQA, HellaSwag, and MMLU.
@@
Task Metric LLaMA-3.2-1B Full Finetune Unlearn Baseline F-MarI KL-GA GA GD DPO
ARC-Easy acc 0.65 ± 0.01 0.59 ± 0.01 0.60 ± 0.01 0.58 ± 0.01 0.55 ± 0.01 0.42 ± 0.01 0.57 ± 0.01 0.59 ± 0.01
acc_norm 0.61 ± 0.01 0.56 ± 0.01 0.55 ± 0.01 0.56 ± 0.01 0.53 ± 0.01 0.39 ± 0.01 0.55 ± 0.01 0.56 ± 0.01
ARC-Challenge acc 0.31 ± 0.01 0.30 ± 0.01 0.31 ± 0.01 0.29 ± 0.01 0.30 ± 0.01 0.27 ± 0.01 0.30 ± 0.01 0.30 ± 0.01
acc_norm 0.36 ± 0.01 0.35 ± 0.01 0.36 ± 0.01 0.34 ± 0.01 0.33 ± 0.01 0.29 ± 0.01 0.32 ± 0.01 0.32 ± 0.01
PIQA acc 0.74 ± 0.01 0.71 ± 0.01 0.71 ± 0.01 0.71 ± 0.01 0.70 ± 0.01 0.63 ± 0.01 0.69 ± 0.01 0.71 ± 0.01
acc_norm 0.74 ± 0.01 0.71 ± 0.01 0.71 ± 0.01 0.69 ± 0.01 0.69 ± 0.01 0.62 ± 0.01 0.70 ± 0.01 0.72 ± 0.01
Hellaswag acc 0.48 ± 0.00 0.48 ± 0.00 0.50 ± 0.00 0.47 ± 0.00 0.46 ± 0.00 0.30 ± 0.00 0.45 ± 0.00 0.47 ± 0.00
acc_norm 0.64 ± 0.00 0.63 ± 0.00 0.65 ± 0.00 0.62 ± 0.00 0.60 ± 0.00 0.38 ± 0.00 0.59 ± 0.00 0.61 ± 0.00
MMLU acc 0.37 ± 0.00 0.30 ± 0.00 0.31 ± 0.00 0.28 ± 0.00 0.26 ± 0.00 0.23 ± 0.00 0.26 ± 0.00 0.28 ± 0.00
- humanities acc 0.35 ± 0.01 0.30 ± 0.01 0.31 ± 0.01 0.29 ± 0.01 0.26 ± 0.01 0.24 ± 0.01 0.27 ± 0.01 0.29 ± 0.01
- other acc 0.41 ± 0.01 0.31 ± 0.01 0.32 ± 0.01 0.29 ± 0.01 0.28 ± 0.01 0.24 ± 0.01 0.25 ± 0.01 0.30 ± 0.01
- social sciences acc 0.39 ± 0.01 0.32 ± 0.01 0.32 ± 0.01 0.28 ± 0.01 0.26 ± 0.01 0.22 ± 0.01 0.24 ± 0.01 0.26 ± 0.01
- stem acc 0.33 ± 0.01 0.26 ± 0.01 0.28 ± 0.01 0.28 ± 0.01 0.25 ± 0.01 0.21 ± 0.01 0.27 ± 0.01 0.26 ± 0.01
Table 6: Comprehensive evaluation results across multiple benchmarks for the LLaMA-3.2-1B unlearning experiment.
@@ E.2 Multiple Detection Test Results
Here, we provide more ablation details on the undetectability result provided in Figure 5, Section 4.4.
(a) Unlearn Baseline
(b) Baseline
(c) Forgetting-MarI